jhonyy9 (jhonyy9):

why can being so difficile to solve this ? 9 - 3/(1/3) +1 = ?

jhonyy9 (jhonyy9):

my opinion that is equal 1

OpenStudy (abdullahm):

I agree with your answer.

jhonyy9 (jhonyy9):

but do you know the way how you get this result of one ?

OpenStudy (lowkey.s):

1

OpenStudy (abdullahm):

When we use PEMDAS or BODMAS We need to do the multiplication or division first, from left to right before we do the addition or subtraction from left to right 9 - 3/(1/3) +1 = 9 - 9 + 1 = 0 + 1 = 1

OpenStudy (lowkey.s):

yep!

OpenStudy (mathmale):

I view this differently. 9 - 3/(1/3) +1 can be re-written as 9 - (3/3) + 1 = 9. End.

OpenStudy (abdullahm):

Mathmale, I have one issue with your reply. How does 3/(1/3) = 3/3? 3/(1/3) = 3 * 3/1 = 9 Clearly, 3/(1/3) cannot be 3/3

OpenStudy (otherworldly):

He rewrote it theres a different way of solving this problem

OpenStudy (mathmale):

Abdullah: You're correct, of course. Thanks for pointing out my error.